• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name J. Singh

    Log entry time 04:34:09 on October19,2005

    Entry number 157376

    keyword=happex → spectrometer → happex


    Rob is here and requests that we switch from the HAPPEX DAQ to the Spectrometer DAQ so that he can do an alignment check. This was recommended by Bob b/c the right HRS has been down for the last few days.

    0012 checked operation of Spec DAQ by running spot++, run 2872 0014 called MCC and requested beam current of 1 µA, got 1.2 0016 recorded and changed HAPPEX detector HV left old (set V) new (set V) 5th PR, ch 8 -1390 -1900 5th PR, ch 9 -1490 -1900 right old (set V) new (set V) S2 1-10 Left, ch 10 -1395 -1900 S2 1-10 Left, ch 11 -1465 -1900 0020 switched cable from HAPPEX → HRS 0022 checked beam position at 1.2 µA, Rob says they look OK 0024 turned on HV for scintillators, S0 and S2 only for both arms 0026 checked xscaler @ 1.2 µA 0028 request beam current to 0.2 µA, got 0.24 xscaler 1.194 µA; 0.241 µA ρ = 4.95 T1 686 kHz 151 kHz 4.54 T2 556 kHz 117 kHz 4.75 T3 648 kHz 142 kHz 4.56 T4 554 kHz 118 kHz 4.69 0032 set VDCs to 4.0 kV, automatically ramping up at about 1kV/min 0037 VDCs reach 4.0 kV 0039 start SPEC DAQ run 2873 0051 3873 ended, backing out of SPEC DAQ mode 0100 VDC's off 0102 beam current spiked to 11.1 µA for about 30 sec, see HALOG 15735
    0110 Rob takes a few SPEC DAQ runs to verify integrity of detectors
    0116 S2/S0 HV off, left and right top 0117 checked other detector HV settings, Rob unplugged cable 0119 resetting HAPPEX detectors HV left old (set V) new (set V)
    5th PR, ch 8 -1900 -1390 5th PR, ch 9 -1900 -1490

    right old (set V) new (set V)
    S2 1-10 Left, ch 10 -1900 -1395
    S2 1-10 Left, ch 11 -1900 -1465

    0122 ask for position feedback off and 10 µA beam to see if the beam drifted before and after the alignment test

    time 00:29 0102 0127
    BPM before (mm) during spike (mm) after (mm)
    1H04A.X +1.55 +1.60 +1.60
    1H04A.Y +0.10 +0.07 +0.10
    1H04B.X -1.41 -1.40 -1.37
    1H04B.y +0.26 +0.21 to +0.39 +0.40

    0133 ask for 58 µA with position feedback on

    figure 1. beam position strip
    figure 2. HV for Left and Right Top



    FIGURE 1

    FIGURE 2